# Constrained least squares simplicial-simplicial regression

### v1

@@ 1 Introduction
An important issue with compositional data analysis is the presence of zeros that prohibit the use of the logarithmic transformations, and hence the approach of Hron et al. (2012), Wang et al. (2013) and Chen et al. (2017), an issue that is not addressed in most papers. The classical strategy addressing this issue is to replace the zero values by a small quantity (Aitchison, 2003). However, the approach of Alenazi (2019) and handles the zero cases in a natural manner. This is not true in general for the TFLR model though.
@@ 3.1 The SCLS model and quadratic programming
The ${\bf A}^{\top}$ matrix can be broken down two three sub-matrices, the $D_{r}\times D_{rp}$ ${\bf A}_{1}$, the $D_{rp}\times D_{rp}$ ${\bf A}_{2}$ and the the $D_{rp}\times D_{rp}$ ${\bf A}_{3}$, one stacked under the other,
@@
  • The ${\bf A}_{1}^{\top}$ matrix contains the identity matrix ${\bf I}_{D_{r}}$, $D_{p}$ times, one next to the other
    ${\bf A}_{1}^{\top}=\left({\bf I}_{D_{r}},\ldots,{\bf I}_{D_{r}}\right).$
@@
    ${\bf A}_{1}^{\top}{\bf b}=\left({\bf I}_{D_{r}},\ldots,{\bf I}_{D_{r}}\right){\bf b}=\left(1,\ldots,1\right)^{\top}.$
@@ 3.6 More than one simplicial predictors
where $a_{m}\geq 0$ for $m=1,\ldots,M$ and $\sum_{m=1}^{M}=1$, are the weights assigned to each set of simplicial predictors. Thus, the previous case (27) can be seen as a special case of (28) where $a_{1}=\ldots=a_{m}=1/M$. The weights may be computed using quadratic programming again, but in doing so the dimensionality of the required matrix $\bf D$ will explode as the number of simplicial predictors and the number of components of each increase, thus one can use simpler optimizers in R. The possibility of allowing for more than one simplicial predictors in either model opens the door to conditional association testing and subsequently to simplicial variable selection.

### v2

@@ 1 Introduction
An important issue with compositional data analysis is the presence of zeros that prohibit the use of the logarithmic transformations, hence the approach of Hron et al. (2012), Wang et al. (2013) and Chen et al. (2017) cannot be applied, an issue that is not addressed in most papers. The classical strategy addressing this issue is to replace the zero values by a small quantity (Aitchison, 2003). However, the approach of Alenazi (2019) handles the zero cases in a natural manner. This is not true in general for the TFLR model though.
@@ 3.1 The SCLS model and quadratic programming
The ${\bf A}^{\top}$ matrix can be broken down two three sub-matrices, the $D_{p}\times D_{rp}$ ${\bf A}_{1}$, the $D_{rp}\times D_{rp}$ ${\bf A}_{2}$ and the the $D_{rp}\times D_{rp}$ ${\bf A}_{3}$, one stacked under the other,
@@
  • The ${\bf A}_{1}^{\top}$ matrix contains the identity matrix ${\bf I}_{D_{p}}$, $D_{r}$ times, one next to the other
    ${\bf A}_{1}^{\top}=\left({\bf I}_{D_{p}},\ldots,{\bf I}_{D_{p}}\right).$
@@
    ${\bf A}_{1}^{\top}{\bf b}=\left({\bf I}_{D_{p}},\ldots,{\bf I}_{D_{p}}\right){\bf b}=\left(1,\ldots,1\right)^{\top}.$
@@ 3.6 More than one simplicial predictors
where $a_{m}\geq 0$ for $m=1,\ldots,M$ and $\sum_{m=1}^{M}a_{m}=1$, are the weights assigned to each set of simplicial predictors. Thus, the previous case (27) can be seen as a special case of (28) where $a_{1}=\ldots=a_{m}=1/M$. The weights may be computed using quadratic programming again, but in doing so the dimensionality of the required matrix $\bf D$ will explode as the number of simplicial predictors and the number of components of each increase, thus one can use simpler optimizers in R. The possibility of allowing for more than one simplicial predictors in either model opens the door to conditional association testing and subsequently to simplicial variable selection.
@@ 3.8.2 Categorical predictor or response
3.9 Alternative implementation of the TFLR model and generalizations thereof
The $\bf D$ matrix in Eq. (19) is actually the cross-product of the following design matrix that contains the simplicial predictor
${\bf X}^{*}={\bf I}_{D_{r}}\otimes{\bf X}=\left(\begin{array}[]{cccc}{\bf X}&{\bf 0}&\ldots&{\bf 0}\\ {\bf 0}&\ddots&{\bf 0}&\vdots\\ \vdots&{\bf 0}&\ddots&{\bf 0}\\ {\bf 0}&\ldots&{\bf 0}&{\bf X}\end{array}\right).$ (34)
The simplicial response can be vectorized ${\bf y}^{*}=\mathop{\operator@font vec}\nolimits({\bf Y})$ and then estimate the coefficients of a logistic regression model of ${\bf y}^{*}$ on ${\bf X}^{*}$ with the identity as the link function, no constant term and imposing the same constraints as in the SCLS. This strategy is the same as the in one used in the seemingly unrelated regression equations (SURE) model (Zellner, 1962). The SURE idea can also be used if one wants to substitute the square loss (13), or the KLD loss (12) with the $L_{1}$ norm loss and thus perform a simplicial constrained multivariate median regression (or spatial median regression (Chakraborty, 2003)). The median regression has been implemented in the quantreg R package (Koenker, 2023) and that implementation allows for inequality constraints and hence the constraints will be nearly the same to those used in Eq. (14).